OpenStudy (anonymous):

Suppose f and g are functions differentiable on (0, +infinity). If \[f(2\sqrt{x})=x^2g(\sqrt{x}+2)\] and it is known that f'(4)=7 and g'(4)=-3, determine g(4).

zepdrix (zepdrix):

Hey there :) Differentiate each side with respect to x. You'll need to do some product rule business on the right side, and then a bunch of chain rule after that, ya? Since you're applying product rule, you should end up with both g and g', so you'll be able to solve for g with some algebra.

OpenStudy (anonymous):

Hi! And thank you! :)

zepdrix (zepdrix):

\[\large\rm f(2\sqrt{x})=x^2g(\sqrt{x}+2)\]Differentiating,\[\large\rm f'(2\sqrt{x})\cdot\color{royalblue}{(2\sqrt x)'}=\color{royalblue}{(x^2)'}g(\sqrt{x}+2)+x^2\color{royalblue}{(g(\sqrt{x}+2))'}\]So what I've done here is, Applied chain rule to the left side, and only "set up" the product rule on the right so far.

zepdrix (zepdrix):

Confusing? :o What do you think?

OpenStudy (anonymous):

It's not confusing so far. I think I can figure out the rest.

zepdrix (zepdrix):

cool :)